OpenStudy (anonymous):

what is 9ft 11in - 4ft 6in

OpenStudy (anonymous):

5 ft 5 inches!

OpenStudy (anonymous):

Well, I can see where you would get that. But I think you may be wrong.

OpenStudy (anonymous):

Really? O.o

OpenStudy (anonymous):

Yeah. I think we have to go like: 9*12=108+11=119. And then 4*12=48+6=54. 119-54=65. Now we have to find out what it would be with feet and inches. 5 feet 5 inches... O_O oh wait. yeah you are correct sorry XD.